# Carbon-Aware Computing for Data Centers with Probabilistic Performance Guarantees

### v2

@@ III Distributionally-robust scheduling
^$\hat{\mathbb{P}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{s^{i}}\ ,$ (10)
@@
$\mathcal{B}^{\varepsilon}:=\left\{\mathbb{Q}\in\mathcal{P}_{1}(\mathcal{S})\leavevmode\nobreak\ |\leavevmode\nobreak\ d_{W}\left(\hat{\mathbb{P}},\mathbb{Q}\right)\leq\varepsilon\right\}\ ,$ (11)
where $\mathcal{P}_{1}(\mathcal{S})$ is the set of Borel probability measures with finite first moment, and $d_{W}\big{(}\hat{\mathbb{P}},\mathbb{Q}\big{)}$ is the so-called Wasserstein distance, given in Definition 2, between the probability distributions $\mathbb{Q}$ and $\hat{\mathbb{P}}$ [33]. Intuitively, (11) describes the set of distributions that are within a radius $\varepsilon$, as measured by the Wasserstein metric $d_{W}$ of the empirical distribution $\hat{\mathbb{P}}$. Ambiguity sets based on the Wasserstein distance are expressive and particularly well-suited for modeling and robustifying against so-called black-swan events, namely, rare and unpredictable outlier events with extreme impact [^3] For instance, an extremely large compute load that, under classical placement schemes, would lead to execution delays for time-sensitive jobs. [34].
@@
This approach enables us to achieve robust (probabilistic) constraint satisfaction while providing a probabilistic guarantee on the realized cost.
Remarkably, the worst-case constraint over the set of distribution in the ambiguity set that appears in problem (12) admits a tractable reformulation as a linear program (LP) that depends on the observed samples $s^{i}\in\mathcal{S}$ with $\mathcal{S}$ being an a priori known support set for $\mathbb{P}$ defined here as
$\mathcal{S}:=\{s\in\mathbb{R}^{KC}\leavevmode\nobreak\ |\leavevmode\nobreak\ Gs\leq h\}.$ (13)
@@
IV Real-time job placement
@@
We point out that in the DRO problem used for day-ahead planning it is assumed that the continuous aggregate load profile can be split into arbitrary small fractions given by the scheduling strategy $Y_{k,c,t,d}$. However, in reality the aggregate load $s_{k,c}$ of class $c$ at time $k$ is the sum of compute volumes of discrete jobs $s_{k,c}=\sum_{j}u_{c}^{j,\text{vol}}$. Thus, in practice, the extent to which the optimal scheduling strategy $Y^{*}$ can be approximated with discrete jobs depends on their size, number and runtime. More details on application specific considerations and extensions are given in Subsection V-A.
In real-time operation, a job $u$ of class $c$ arrives, and we want to place it in cluster $d$ and time slot $t$ which is furthest away from fulfilling the optimal load fraction $Y^{*}$. Thus, for ${}^{\text{th}}$ job of volume $u_{c}^{j,\text{vol}}$ arriving in hour $k$, we choose the placement tuple $(d,t)$ for which the difference between $Y^{*}$ and the current load ratios $\frac{Y^{\text{abs}}_{k,c,t,d}}{\sum_{i=0}^{j-1}u_{c}^{i,\text{vol}}}$ is maximal, expressed in the following:
@@
V Implementation and Numerics
V-A Application specific considerations and extensions
@@
    Job runtimes: We currently assume job runtimes of one time step. To handle longer runtimes, we can introduce additional job classes and constraints to ensure time continuity of the scheduling strategy.
@@
    Capacity constraints: The VCCs $v_{t,d}$ sent to clusters can be used in two ways: (i) As information for the cluster-level operating system (known as Borg at Google [37, 36]) on the CVAR bound of the daily load profile; (ii) As a hard capacity constraint for the placement algorithm. The implementation by Google in [26] incorporates both (i) and (ii). We choose approach (i) for the simulations in Section V.B-F as it aligns with the theoretical derivations in Section II-III. For Section V.G we use approach (ii) which is closer to the practical implementation.
@@
V-B Simulation set-up and processing of data
@@
V-C Day-ahead: Training and validation scenarios
@@
V-D Greedy vs planning
In this study, we compare the optimal scheduling strategy resulting from our proposed DRO approach with a basic greedy approach that places every incoming job at hour $k$ to the “cheapest” cluster within the job classes’ spatial range $d\in\mathcal{D}_{c}$ and with available (cluster) machine capacity. The load profiles over one day for both policies are shown in Figure 4. Table II shows the cost across all 15 validation scenarios compared to an ideal placement that has a perfect forecast of the next day’s aggregate load $s_{k,c}$.
@@ V-D Greedy vs planning
Table II: Comparing the cost of the DRO and greedy policy to the optimal cost with perfect forecast.
@@
  • The costs of the DRO strategy are just 2.57 % higher than those of the policy with perfect forecast for $\varepsilon=8\cdot 10^{-3}$ and 8.86 % higher for $\varepsilon=5\cdot 10^{-2}$. For both robustness levels, it significantly reduces cost compared to the greedy policy, as shown in Table II.
@@
Figure 4: Comparing the running load of the DRO and greedy policy for discrete jobs over one day.
V-E Exploiting spatial range and temporal flexibility
@@
In this study, we aim to demonstrate that the DRO scheduling strategy utilizes the spatial range and temporal flexibility of jobs to shift them to less expensive hours and clusters. In Figure 5, we present the optimal load fractions for all spatially flexible classes: DC-flexible classes $c=4$ and $c=5$, and the globally flexible class $c=6$. The load distribution for these classes aligns with cluster-level price curves to avoid high-cost periods. For example, in the globally flexible class $c=7$, all load shifts to $d=1$ in the early hours to minimize costs.
@@
In Figure 6, we illustrate how the DRO scheme exploits temporal flexibility with the two time axes, submission and execution time. Without the second time axis (i.e., if $t=k$), the heatmap would show only diagonal entries, indicating 100% execution at submission time, thereby eliminating any temporal flexibility and enforcing that all loads are processed immediately. Figure 6 also highlights the effect of the tuning parameter $\beta$. For $\beta=0.05$, much of the load is executed immediately upon submission to meet the delay constraint of $h_{c}=10$ hours. In contrast, for $\beta=0.7$, the probabilistic constraint in (6) is relaxed, allowing compute loads to be pushed to later hours to minimize costs.
@@
V-F Influence of CVaR constraint parameters $\beta$ and $\varepsilon$
(a) With $\beta=0.02$ and 0 VCC violations.
(b) With $\beta=0.1$ and 4 VCC violations.
(c) With $\beta=0.5$ and 33 VCC violations.
In this section, we highlight the role of the CVaR level $\beta$ and DRO radius $\varepsilon$ influencing the CVaR constraint in (6). Figure 7 shows the load profiles of discrete jobs over one day for increasing values of the CVaR level $\beta$. We observe that VCCs $v_{d,t}$ tighten with increasing $\beta$, corresponding to a relaxation of the probabilistic constraint against violating the VCC limits. While a value of $\beta=0.5$ allows for significant shifting of compute load in time, thereby reducing operational costs, it may result in (cluster) level machine capacity being violated and some jobs not being executed due to insufficient capacity. However, note that in practice, the DRO would be implemented in a receding-horizon fashion, in which case it is not necessary for all jobs to be executed by the end of the finite horizon.
Figure 8 shows how $\varepsilon$, the ambiguity set radius in (11), is used as a tuning knob to trade-off between performance and robustness against future realizations of the compute loads. The problem reduces to SAA with $\varepsilon=0$. A choice of $\varepsilon=10^{-2}$ achieves a good balance between performance and VCC constraint violations across all 15 validation scenarios.
V-G Practical implementation
In the real application presented by Google in [26], VCC curves serve as a hard constraint for resource availability and the cluster-level operating system. When the load pushed to a specific cluster exceeds its planned capacity, jobs are placed in a cluster-level queue and executed as resources become available. In this numerical study, we tested how our proposed real-time job placement scheme in (15) handles discrete jobs. Jobs have a one-hour runtime and volume size drawn from a per-class normal distribution $\mathcal{N}_{c}(u_{c}^{\text{vol,mean}},0.1)$, with $u_{c}^{\text{vol,mean}}$ being the per-class compute load divided by job count.
@@ V-G Practical implementation
VI Conclusion

### v3

@@ IV Distributionally-robust scheduling
$\hat{\mathbb{P}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{s^{i}}\ ,$ (15)
@@
$\mathcal{B}^{\varepsilon}:=\left\{\mathbb{Q}\in\mathcal{P}_{1}(\mathcal{S})\penalty 10000\ |\penalty 10000\ d_{W}\left(\hat{\mathbb{P}},\mathbb{Q}\right)\leq\varepsilon\right\}\ ,$ (16)
[…]
@@
This approach enables us to achieve robust (probabilistic) constraint satisfaction while providing a probabilistic guarantee on the realized cost. Furthermore, the coupled optimization of VCCs and scheduling $Y$, and the use of an ambiguity set centered on the empirical distribution of the available data distribution, allow us to exploit the temporal and spatial correlations that exist across job classes.
Remarkably, the worst-case constraint over the set of distribution in the ambiguity set that appears in problem (17) admits a tractable reformulation as a linear program (LP) that depends on the observed samples $s^{i}\in\mathcal{S}$, with $\mathcal{S}$ being an a priori known support set for $\mathbb{P}$ defined here as
$\mathcal{S}:=\{s\in\mathbb{R}^{KC}\penalty 10000\ |\penalty 10000\ Gs\leq h\}.$ (18)
@@
The number of decision variables scales quadratically with $T$, linearly with $C$ and $D$. Further, the formulation naturally lends itself to decomposition techniques based on spatial constraints of jobs.
@@
The proposed method differs from the approach in [30] in the formulation of the constraints and in how robustness is introduced in the optimization problem. In [30], robustness is obtained using the $97^{th}$ percentile bound of the forecasted compute load. In contrast, we achieve robustness by formulating the DRO (13). The level of robustness is controlled by the Wasserstein ambiguity set radius $\varepsilon$, while the risk level can be adjusted by varying the $\beta$ parameter of the CVaR-based constraint. Compared to the approach in [30], our proposed co-design of VCCs and scheduling enables directly exploiting the correlation of load profiles between classes which emerges from the data, as captured by the empirical compute load distribution $\hat{\mathbb{P}}$.
V Real-time job placement
@@
We point out that in the DRO problem used for day-ahead planning it is assumed that the continuous aggregate load profile can be split into arbitrary small fractions given by the scheduling strategy $Y_{k,c,t,d}$. However, in reality the aggregate load $s_{k,c}$ of class $c$ at time $k$ is the sum of compute volumes of discrete jobs $s_{k,c}=\sum_{j}u_{c}^{j,\text{vol}}$. Thus, in practice, the extent to which the optimal scheduling strategy $Y^{*}$ can be approximated with discrete jobs depends on their size, number and runtime. In case of unexpected real-time changes, such as machine failure, the VCC limits will be recomputed the next day reducing the affected cluster’s capacity and redistributing load. If one would like the approach to handle cluster or machine failures in real-time, then a protection mechanism could be implemented in the job placement algorithm.
More details on application specific considerations and extensions are given in Subsection VI-A.
In real-time operation, a job $u$ of class $c$ arrives, and we want to place it in cluster $d$ and time slot $t$ which is furthest away from fulfilling the optimal load fraction $Y^{*}$. Thus, for the $^{\text{th}} j$ob of volume $u_{c}^{j,\text{vol}}$ arriving in hour $k$, we choose the placement tuple $(d,t)$ for which the difference between $Y^{*}$ and the current load ratios $\frac{Y^{\text{abs}}_{k,c,t,d}}{\sum_{i=0}^{j-1}u_{c}^{i,\text{vol}}}$ is maximal, expressed in the following:
@@
The computational complexity of our real-time job placement scheme (20) is determined by finding the maximum value in a vector/tensor. This operation has a complexity of $O(D\texttimes C)$, where $D$ is the number of clusters, and $C$ is the number of job classes.
VI Implementation and Numerics
In this section, we discuss relevant extensions to our model tailored to real-world application and present illustrative simulation studies to showcase the main features and principles of our DRO scheduling strategy. To keep the focus clear and illustrative, we consider sinusoidal cost function parameters as a representative example of time-varying grid signals. Further, we point out that the simulations are presented for 24hrs of execution time instead of $T=K+\max h_{c}=34$hrs as new VCC limits get pushed every day. We assume that remaining jobs will fit within the VCC limits of the next day which is consistent with the approach in [30].
VI-A Application specific considerations and extensions
@@
    Job runtimes: We currently assume homogeneous job runtimes of one time step. This is consistent with the industry practice of splitting large compute jobs into small parallelizable tasks. To handle longer runtimes, we can introduce additional job classes and constraints to ensure time continuity of the scheduling strategy.
@@
    Capacity constraints: The VCCs $v_{t,d}$ sent to clusters can be used in two ways: (i) As information for the cluster-level operating system (known as Borg at Google [51, 50]) on the CVaR bound of the daily load profile; (ii) As a hard capacity constraint for the placement algorithm. The implementation by Google in [30] incorporates both (i) and (ii). We choose approach (i) In Section VI-B through Section VI-E as it aligns with the theoretical derivations in Section III and IV. For Section VII we use approach (ii) which is closer to the practical implementation.
@@
VI-B Simulation set-up and processing of data
@@
VI-C Day-ahead: Training and validation scenarios
@@
VI-D Greedy vs planning
In this study, we compare the optimal scheduling strategy resulting from our proposed DRO approach with a basic greedy approach that places every incoming job at hour $k$ to the “cheapest” cluster within the job classes’ spatial range $d\in\mathcal{D}_{c}$ and with available (cluster) machine capacity. The load profiles over one day for both policies are shown in Figure 4. Table III shows the cost across all 15 validation scenarios compared to an ideal placement that has a perfect forecast of the next day’s aggregate load $s_{k,c}$.
Table III: Comparing the cost of the DRO and greedy policy to the optimal cost with perfect forecast.
@@ VI-D Greedy vs planning
  • The costs of the DRO strategy are just 2.57 % higher than those of the policy with perfect forecast for $\varepsilon=8\cdot 10^{-3}$ and 8.86 % higher for $\varepsilon=5\cdot 10^{-2}$. For both robustness levels, it significantly reduces cost compared to the greedy policy, as shown in Table III.
@@
Figure 4: Comparing the running load of the DRO and greedy policy over one day.
VI-E Influence of CVaR constraint parameters $\beta$ and $\varepsilon$
(a) With $\beta=0.02$ and 0 VCC violations.
(b) With $\beta=0.1$ and 4 VCC violations.
(c) With $\beta=0.5$ and $3$ VCC violations.
In this section, we highlight the role of the CVaR level $\beta$ and DRO radius $\varepsilon$ influencing the CVaR constraint in (7). The parameter $\beta$ is a design parameter chosen to control the risk of constraint violations. By adapting the value of $\beta\in(0,1]$, we can model a continuous spectrum of formulations. Choosing $\beta=1$ results in a problem that requires the constraint to hold in expectation (with respect to the worst-case distribution in the ambiguity set), while choosing $\beta$ near zero results in a constraint that approaches a worst-case robust solution (with respect to the worst-case distribution in the ambiguity set). This parameter should be chosen according to the risk sensitivity of the specific application. Figure 5 shows the load profiles of discrete jobs over one day for increasing values of the CVaR level $\beta$. We observe that VCCs $v_{d,t}$ tighten with increasing $\beta$, corresponding to a relaxation of the probabilistic constraint against violating the VCC limits. While a value of $\beta=0.5$ allows for significant shifting of compute load in time, thereby reducing operational costs, it may result in (cluster) level machine capacity being violated and some jobs not being executed due to insufficient capacity. However, note that in practice, the DRO would be implemented in a receding-horizon fashion, in which case it is not necessary for all jobs to be executed by the end of the finite horizon.
Moreover, it is crucial to select an appropriate radius $\varepsilon$ to guarantee that the true (unknown) distribution $\mathbb{P}$ belongs to the ambiguity set $\mathcal{B}^{\varepsilon}$ in (16). When samples used to construct the ambiguity set center $\hat{\mathbb{P}}$ are drawn from some stationary $\mathbb{P}$, concentration results can establish confidence bounds on the Wasserstein distance between $\hat{\mathbb{P}}$ and $\mathbb{P}$ [34], and thus provide guidance on how to choose the radius $\varepsilon$. However, in practical applications, these theoretical bounds often prove overly conservative. Moreover, the assumption of a stationary load-generating distribution may not hold in many real-world scenarios. The parameter $\varepsilon$ is, in practice, treated as an empirical tuning knob, at the expense of formal theoretical guarantees. A practical way of picking a value of the radius that is not too conservative is to perform a leave-one-out cross validation, using N1N-1 scenarios to build the empirical distribution $\hat{\mathbb{P}}$ and test against the left-out sample for different values of $\varepsilon$. The choice of radius can be guided by selecting a value of $\varepsilon$ that balances cost and constraint violations. Figure 6 shows how different values of $\varepsilon$, provide different trade-offs between performance and robustness against previously unseen future realizations of the compute loads. The problem reduces to SAA with $\varepsilon=0$ and a robust optimization problem for $\varepsilon$ large. A choice of $\varepsilon=10^{-2}$ achieves a good balance between performance and VCC constraint violations across all 15 validation scenarios.
VI-F Exploiting spatial range and temporal flexibility
@@
In this study, we aim to demonstrate that the DRO scheduling strategy utilizes the spatial range and temporal flexibility of jobs to shift them to less expensive hours and clusters. In Figure 7, we present the optimal load fractions for all spatially flexible classes: DC-flexible classes $c=4$ and $c=5$, and the globally flexible class $c=6$. The load distribution for these classes aligns with cluster-level price curves to avoid high-cost periods. For example, in the globally flexible class $c=7$, all load shifts to $d=4$ in the early hours to minimize costs.
@@
In Figure 8, we illustrate how the DRO scheme exploits temporal flexibility with the two time axes, submission and execution time. Without the second time axis (i.e., if $t=k$), the heatmap would show only diagonal entries, indicating 100% execution at submission time, thereby eliminating any temporal flexibility and enforcing that all loads are processed immediately. Figure 8 also highlights the effect of the tuning parameter $\beta$. For $\beta=0.02$, much of the load is executed immediately upon submission to meet the delay constraint of $h_{c}=10$ hours. In contrast, for $\beta=0.2$, the probabilistic constraint in (7) is relaxed, allowing compute loads to be pushed to later hours to minimize costs.
@@
VII Realistic implementation with carbon cost and power constraints
In the real application presented by Google in [30], VCC curves serve as a hard constraint for resource availability and the cluster-level operating system. When the load pushed to a specific cluster exceeds its planned capacity, jobs are placed in a cluster-level queue and executed as resources become available. Note that our approach is fully agnostic to the low level control of the cluster, in the real system placement to machines in a cluster is handled by a system called Borg [50]. Here, we choose to implement a simple “First-In-First-Out” queuing system. In this numerical study, we tested how our proposed real-time job placement scheme in (20) handles discrete jobs. Jobs have a one-hour runtime and volume size drawn from a per-class normal distribution $\mathcal{N}_{c}(u_{c}^{\text{vol,mean}},0.1)$, with $u_{c}^{\text{vol,mean}}$ being the per-class compute load divided by job count.
VII-A Queue length increases with tighter VCCs
@@
VII-B Ramping limits affect VVCs and cost
[…]
Figure 10: Showing load profiles (black) for 15 validation scenarios $s^{i}_{\text{val}},i\in\mathbb{Z}_{15}$ and capacity curves VCCRL,d (blue tones) under ramp constraints as defined in (9), and comparing them to VCC limits without ramp constraints (magenta).
VII-C Cost function elements and VCC effects
Figure 11: Comparing load profiles using real carbon data sourced from[41] for four distinct Google DC locations.
We use real carbon data for four distinct locations of Google DCs: (i) Berkeley, South Carolina, (ii) Henderson, Nevada, (iii) Saint-Ghislain, Belgium, (iv), Quilicura, Chile. We use the carbon intensity data from Electricity Maps [41], which is the official company that Google sources its carbon data from in real operation. For the presented simulation we used data from the 22nd of February, starting at 10 am. The results are presented in Figure 11. First of all, we notice that the scheduling strategy exploits the geographic flexibility of compute load to decrease the load in the carbon intense location of Berkeley in South Carolina. Only load that is spatially inflexible is placed here. In addition, we notice that the scheduler also exploits the temporal flexibility of jobs in Henderson, Nevada where carbon intensity greatly drops from hour 8 to 12 am.
VIII Conclusion